OpenStudy (jack1):

cost = P = (20 + 0.5x) + 0.15(20 + 0.5x)

OpenStudy (jack1):

P = (20 + 0.5x) + 0.15(20 + 0.5x) cost = 1.15 (20 +0.5x) so if you have 62.10 62.10 = 1.15 (20 + 0.5x) 62.10 /1.15 = (20 + 0.5 x) can u solve from here @samantha429 ?

OpenStudy (anonymous):

yes thank you

OpenStudy (jack1):

all good ;)